### v1

# Zhuosheng Zhang♣,†, Gongshen Liu♣,†

Abstract
Autonomous graphical user interface (GUI) agents powered by multimodal large language models have shown great promise. However, a critical yet underexplored issue persists: over-execution, where the agent executes tasks in a fully autonomous way, without adequate assessment of its action confidence to compromise an adaptive human-agent collaboration. This poses substantial risks in complex scenarios, such as those involving ambiguous user instructions, unexpected interruptions, and environmental hijacks. To address the issue, we introduce OS-Kairos, an adaptive GUI agent capable of predicting confidence levels at each interaction step and efficiently deciding whether to act autonomously or seek human intervention. OS-Kairos is developed through two key mechanisms: (i) collaborative probing that annotates confidence scores at each interaction step; (ii) confidence-driven interaction that leverages these confidence scores to elicit the ability of adaptive interaction. Experimental results show that OS-Kairos substantially outperforms existing models on our curated dataset featuring complex scenarios, as well as on established benchmarks such as AITZ and Meta-GUI, with 24.59$\sim$87.29improvements in task success rate. OS-Kairos facilitates an adaptive human-agent collaboration, prioritizing effectiveness, generality, scalability, and efficiency for real-world GUI interaction. The dataset and codes are available at https://github.com/Wuzheng02/OS-Kairos.
[Uncaptioned image] OS-Kairos: Adaptive Interaction for MLLM-Powered GUI Agents
Pengzhou Cheng♣,thanks:   Equal contribution. Corresponding authors., Zheng Wu♣,∗, Zongru Wu, Aston Zhang, Zhuosheng Zhang♣,†, Gongshen Liu♣,† School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University {cpztsm520,wzh815918208,wuzongru,zhangzs,lgshen}@sjtu.edu.cn
1 Introduction
Figure 1: Illustration of GUI agents executing a complex shopping instruction across two paradigms: (a) Autonomous, where the agent cannot complete the task independently; (b) Adaptive, where the agent dynamically adjusts its autonomy based on confidence levels.
Multimodal large language models (MLLMs) have been explored to develop graphical user interface (GUI) agents capable of analyzing the screen and performing human-like behaviors on operating systems Hong et al. (2024); Zhang et al. (2024a); Wang et al. (2024a). Existing efforts in building GUI agents have focused on the autonomous mode Niu et al. (2024); Zhang et al. (2024b); Nguyen et al. (2024); Liu et al. (2024), with improved capabilities such as grounding (Wu et al., 2024b; Qin et al., 2025) and reasoning (Zhang and Zhang, 2024; Zhang et al., 2024b; Liu et al., 2025). Despite exciting progress, we observe that existing GUI agents exhibit significant over-execution issues — the agent executes tasks in a fully autonomous way, without adequate assessment of its action confidence to compromise an adaptive human-agent collaboration. As shown in Figure 1(a), popular GUI agents such as OS-Atlas (Wu et al., 2024b) are unable to click the filters button correctly, causing unexpected interruptions and task failure.
Over-execution poses significant challenges in complex real-world scenarios (Examples shown in Figure 2), highlighting fundamental limitations in current GUI agents. First, ambiguous instructions from the user leads to information absence in GUI automation (e.g., account logout). Second, existing GUI agents depend heavily on the foundation MLLMs and therefore suffer from unexpected interruptions when executing complex instructions. Besides, these models will also generate hallucinations Sridhar et al. (2023) and shortcut predictions Wu et al. (2024b); Zhu et al. (2024). Third, environmental hijacks, such as network connection failure and pop-up hijacking) Ma et al. (2024a).
To address these challenges, we are motivated to integrate confidence scoring into the foundation model, allowing adaptive human intervention for GUI agents (Figure 1(b)). Concretely, we introduce OS-Kairos, an adaptive GUI agent capable of predicting confidence levels at each interaction step and efficiently determining whether to act autonomously or seek human intervention. OS-Kairos incorporates two key mechanisms: (i) collaborative probing that annotates confidence scores at each interaction step; (ii) confidence-driven interaction that utilizes these confidence scores to enhance the ability of adaptive interaction.
To annotate the confidence scores for the probed GUI agents in real-world scenarios, we first design a collaborative confidence probing framework. Inspired by Chen et al. (2024a), this framework integrates a layout parsing model Tang et al. (2019) and the most capable proprietary model, $\mathtt{GPT}$-$\mathtt{4o}$ Achiam et al. (2023) to function as a critic model. The critic model is used to supervise plan scheduling and confidence score based on our curated instructions to address complex scenarios. This framework is the first toolkit designed to identify when human intervention is necessary, generate confidence scores, and facilitate the automated construction of GUI trajectories. To further integrate confidence scoring into the probed GUI agent, we validate and refine these GUI trajectories and then fine-tune the model. This approach ensures action prediction accuracy while improving adaptability of human intervention.
Experimental results in complex scenarios show that OS-Kairos achieves state-of-the-art performance with action type success rate of 99.88%, action success rate of 95.90%, and task success rate of 88.20%. Also, we confirm OS-Kairos’s effectiveness on two well-established GUI benchmarks: Meta-GUI Sun et al. (2022) and AITZ Zhang et al. (2024b). Comprehensive analysis reveals that OS-Kairos prioritizes effectiveness, generality, scalability, and efficiency, making it a competitive agent for real-world GUI interactions. Our work makes the following key contributions:
(i) We introduce OS-Kairos, an adaptive GUI agent that predicts the confidence level of each interaction step and effectively decides whether to act autonomously or seek human intervention.
(ii) We propose a collaborative confidence probing framework for dynamically identifying the confidence scores of the GUI agents in typical complex real-world scenarios, while automatically generating high-quality GUI trajectory.
(iii) We employ confidence-driven interaction to integrate confidence scoring into the GUI agent that forms adaptive human intervention without compromising action prediction.
(iv) We demonstrate that OS-Kairos substantially outperforms existing models on both our curated dataset featuring complex scenarios and well-established benchmarks, with merits of effectiveness, generality, scalability, and efficiency.
2 Related Works
Our work falls into the field of MLLM-powered agents. This section will first review the recent progress in building GUI agents and then discuss the capability probing approaches for GUI agents.
2.1 MLLM-powered GUI Agents
The rise of MLLMs has redefined the paradigm for GUI agents, enabling them to analyze complex screen layouts and generate accurate actions in a more human-like way Zhang et al. (2024a). Importantly, this paradigm is a non-intrusive manner without reliance on complex, platform-specific scripts or predefined workflows. Notable examples across different platforms include SeeAct Zheng et al. (2024) and WebRL Qi et al. (2024) for web navigation, AppAgent Zhang et al. (2023), Auto-UI Zhang and Zhang (2024), and CoCoAgent Ma et al. (2024b) for mobile interactions, and ScreenAgent Niu et al. (2024) for Windows OS applications. This paper investigates the over-execution of MLLM-powered GUI agents on mobile devices.
Early efforts to build GUI agents rely on the availability of commercial MLLMs. These agents can be built through prompt learning based on $\mathtt{GPT}$-$\mathtt{4o}$ or Gemini-Pro Vision, e.g., AppAgent Zhang et al. (2023) and Mobile-Agent Wang et al. (2024a). However, practitioners are concerned about the costs associated with API requests and the delays in inference on mobile devices. Recent studies have focused on fine-tuning to optimize foundation models. On the one hand, they work on performing fine-grained visual understanding Bai et al. (2023), model scaling laws Chen et al. (2024b), multimodal information integration Hong et al. (2024), and GUI grounding enhancements Wu et al. (2024b); Qin et al. (2025) in the pre-training phase. On the other hand, researchers fine-tune the foundation model on GUI-specific datasets to enhance action orientation Wu et al. (2024a), planning decision Zhang et al. (2024c), perception enhancement Ma et al. (2024b), and reasoning Zhang and Zhang (2024); Zhang et al. (2024b). Moreover, a framework based on reinforcement learning (RL) designed specifically for the GUI agents can further enhance robustness Zhou et al. (2024); Liu et al. (2024); Wang et al. (2024b).
Despite the progress, existing GUI agents encounter performance bottlenecks in complex scenarios (Figure 2), such as those involving ambiguous user instructions, unexpected interruptions, and environmental hijacks. Sun et al. (2022) proposed Meta-GUI that leverages precise guidance through task-oriented dialogue. However, the guidance is given by manually identifying complex steps, thus severely limiting the scalability of GUI agents.
2.2 Capability Probing for GUI Agent
GUI agent-oriented capability probing is critical for real-world applications Deka et al. (2017). Generally, the capability of GUI agents can be probed by releasing benchmark datasets. Examples like UIBert Bai et al. (2021), SeeClick Cheng et al. (2024), and OS-Copilot Wu et al. (2024b), which investigate the problem of grounding understanding to UI elements on a screen. Besides, large-scale, diverse, and high-quality trajectory datasets can identify challenges of action prediction in terms of effectiveness (e.g., PixelHelp Li et al. (2020), Meta-GUI Sun et al. (2022), and AndroidWorld Rawles et al. (2024)), task complexity (e.g., Mobile-Bench Deng et al. (2024) and GUI Odyssey Lu et al. (2024)), and data-scaling (e.g., AITW Rawles et al. (2024) and AndroidControl Li et al. (2024)). After identifying the capability bottleneck of GUI agents, the introduction of specific strategies (e.g., planning lists Zhang et al. (2024c), action chains Zhang and Zhang (2024); Zhang et al. (2024b), and supplementary data) further enhance the environment perception. However, most benchmark datasets rely on crowdsourcing and human annotation.
Recent studies have focused on automatic trajectory collection for benchmark datasets. For example, Zhou et al. (2024) introduces a two-stage RL framework that explores successful trajectories during optimization. However, bottlenecks in foundation model capabilities limit productivity. Sun et al. (2024) further proposed OS-Genesis, which back-generates instructions through UI element traversal and ensures the generated high-quality trajectory based on a reward model. However, environment emulators (e.g., Android Studio Emulator Deka et al. (2017)) do not reflect real-world scenarios. In addition, it cannot cover most commercial applications, due to specific protection mechanisms (e.g., RedNote). Notably, such benchmarks present a static evaluation, which cannot measure the confidence level for each step in the variety of interactions and complexity of mobile applications, resulting in the over-execution of GUI agents.
3 Pilot Experiments
In this section, we first define GUI agent paradigms and then investigate the over-execution issue of the existing GUI agent on three complex scenarios. As shown in Figure 2, GUI agents confront substantial risks in real-world scenarios, such as those involving ambiguous user instructions (e.g., information absence and account logout), unexpected interruptions (e.g., hallucinations and shortcuts), and environmental hijacks (e.g., Pop-up hijacking and permission unauthorized).
Figure 2: Illustration of three complex scenarios.
3.1 GUI Agent Paradigm
The task of the GUI agent is defined as a sequence generation problem for MLLMs, with two paradigms: autonomous and interactive.
Autonomous GUI Agent. Given an autonomous GUI agent $\mathcal{F}_{a}$ and system prompt $P$, a user instruction $\tau_{i}\in\mathcal{T}$ can be achieved with continuous interaction steps in the mobile-device environment. At each step $t$, the agent predicts the next action $a_{t}$ followed by $\mathcal{F}_{a}(a_{t}|P(s_{t},h_{t-1},\tau_{i},o_{t}))$, where $s_{t}$ is a screenshot, $h_{t-1}$ is the previous history of the agent $(<s_{1},o_{1},a_{1}>,\cdots,<s_{t-1},o_{t-1},a_{t-1}>)$, and $o_{t}$ is supplementary data (e.g., plan list).
Interactive GUI Agent. Given an interactive GUI agent $\mathcal{F}_{i}$ and system prompt $P$, we expect the agent can be aware of the complex step $t$, and initiate a human intervention. After providing precise guidance $g_{t}$ from a human or advanced model $\mathcal{F}_{s}$, the agent can arrive next step $t+1$, formed by $\mathcal{F}_{s}(a_{t}^{s}|s_{t},h_{t-1},\tau_{i},o_{t})$.
3.2 Challenge of Over-execution
To investigate the performance of existing GUI agents, we randomly select 350 instructions from three complex scenarios (Figure 9 and 10) to evaluate them. Then, we select the autonomous GUI agents: Qwen2-VL-7B and OS-Atlas-Pro-7B, and the interactive GUI agent assisted at each step by $\mathtt{GPT}$-$\mathtt{4o}$ in our pilot evaluation. Following the setting of Zhang and Zhang (2024) and Wu et al. (2024b), we report their performance in terms of action-Type success rate, the step-wise success rate (SR), and task success rate (TSR).
Models Type (%$\uparrow$ SR (%$\uparrow$ TSR (%$\uparrow$
Qwen2-VL-7B 43.19 18.94 0
OS-Atlas-Pro-7B 97.69 59.12 17
Interactive GUI Agent 94.42 86.74 62
Table 1: Pilot evaluation of three complex scenarios. The definition of metrics is deferred to Section 5.1.
Table 1 shows that Qwen2-VL-7B struggles to adapt to complex scenarios, achieving only 43.19% in Type and 18.94% in SR. In contrast, OS-Atlas-Pro-7B, with improved grounding capability, exhibits significant improvement, achieving 97.69% and 59.12% accuracy in Type and SR, respectively. However, the autonomous GUI agents fail to perform effectively on complex steps, resulting in TSR of 0% and 17%. This is attributed to over-execution of the autonomous GUI agent that low SR affects TSR exponentially. In contrast, when using the interactive GUI agent, the SR and TSR can achieve optimal performance, which is enhanced to 86.74% and 62% respectively. However, relying on human intervention for each step is impractical. The effect proof of over-execution and interaction on TSR of GUI agent is further demonstrated in Appendix A.
These observations motivate our exploration of adaptive interaction, where the system can dynamically decide whether to operate autonomously or request human intervention.
4 Methodology
This section presents OS-Kairos. We first introduce a collaborative probing framework that dynamically annotates the confidence scores at each interaction step. Then, we will describe confidence-driven interaction that integrates confidence scoring into GUI agents, resulting in adaptive interaction. Figure 3 shows an overall illustration.
Figure 3: Overall pipeline of OS-Kairos: After collecting instructions for each complex scenario, we annotate confidence scores at each interaction step of the probing agent through a collaborative probing framework. Finally, confidence-driven interaction integrates the adaptive human intervention into the GUI agent, resulting in OS-Kairos.
4.1 Collaborative Probing Framework
This framework integrates instruction collection, confidence annotation, and data refinement, enabling the generation of a high-quality trajectory dataset with a confidence score for each step.
Instruction Collection. We first collect complex instructions $\mathcal{T}=\{\tau_{1},\tau_{2},\cdots,\tau_{N}\}$ from publicly available datasets and human designers, and then expanded by LLMs (e.g., GPT-4) to increase diversity. To comprehensively probe the model’s confidence at each step, these instructions incorporate factors such as language type (both English and Chinese), 12 APPs, and 12 topics. The distribution of APP and topic is shown in Appendix B.
Confidence Annotation. Our confidence probing framework employs an agent-critic collaborative paradigm. To address the challenge of dynamic evaluation and expand coverage to commercial applications, the framework first utilizes Android Studio to connect real mobile devices and establish bidirectional communication with the probed GUI agent $\mathcal{F}_{p}$ deployed at the service station. Second, inspired by Ma et al. (2024b); Wang et al. (2024a), we assume that the layout-parse model enhanced $\mathtt{GPT}$-$\mathtt{4o}$ is the state-of-the-art critic model $\mathcal{F}_{c}$, capable of effectively supervising and guiding $\mathcal{F}_{p}$, thereby ensuring the dynamic probing of the entire trajectory. Additionally, $\mathcal{F}_{c}$ can monitor the entire probing process, including the planned schedule, current step, and instruction completion. The details of the prompt are provided in Appendix C.1.
Specifically, given a user instruction $\tau_{i}$, $\mathcal{F}_{p}$ predicts the next action $a_{t}^{p}$ using the action prompt $P_{p}$ at step $t$, followed by $\mathcal{F}_{p}(a_{t}^{p}|P_{p}(s_{t},h_{t-1},\tau_{i}))$. For example, $\mathcal{F}_{p}$ responds to the instruction $\tau_{i}$ with the first step “Open Amazon APP” as “CLICK <616, 371>”. Meanwhile, $\mathcal{F}_{c}$ first generates a plan schedule $L$ using the prompt $P_{l}$. Based on the current step $L_{t}$ at step $t$, it will also respond with a supervisory action $a_{t}^{c}$ using the action prompt $P_{c}$. Subsequently, $\mathcal{F}_{c}$ evaluates the effectiveness of the current step execution with the scoring prompt $P_{h}$:
$f_{\text{score}}:<\tau_{i},L_{t},s_{t},a_{t}^{p},a_{t}^{c},h_{t-1}>\stackrel{{\scriptstyle\mathcal{F}_{c}}}{{\rightarrow}}\text{score}_{t},$ (1)
where $f_{\text{score}}$ ranges from 1 to 5. When $\text{score}_{t}$ is 5, we consider that $\mathcal{F}_{p}$ is correct to execute the current step, otherwise, the framework will execute action $a_{t}^{c}$ to continue probing the next step until the instruction is judged finished by $\mathcal{F}_{c}$. For instance, $\mathcal{F}_{p}$ provides the action “$\mathtt{SCROLL[UP]}$” at step 5, while the corrective action is “$\mathtt{CLICK<146,357>}$” on the filter button. The framework also incorporates reflective mechanisms to monitor the plan schedule. At each step, $\mathcal{F}_{c}$ determines the completion of instruction $\tau_{i}$ and the current step $L_{t}$:
$f_{t}:\langle L,s_{t}\rangle\xrightarrow{\mathcal{F}_{c}}\text{index},$ (2)
where the framework will retry the current step if $t=\text{index}$, otherwise to continue execution.
Data Refinement. In this phase, we validate and refine these GUI trajectories, ensuring alignment between action and confidence score. The distribution of each step in the complex scenarios is based on its score, as shown in Appendix B. Notably, the distribution of actions scored 5 is concentrated in normal steps, such as “Open APP” or “Click Search Bar”. However, once the instructions contain complex steps, the confidence scores of the probing agent decrease significantly. Hence, we can identify the over-execution steps of the probed GUI agent and treat these steps as requiring advanced GUI agent guidance or human intervention.
4.2 Confidence-driven Interaction
This phase integrates confidence scoring, resulting in a GUI agent with adaptive interaction.
Confidence Scoring Integration. Employing the trajectory from the collaborative probing framework, we introduce OS-Kairos, which integrates confidence scoring with the probed GUI agent $\mathcal{F}_{p}$. Specifically, we employ supervised training to fine-tune $\mathcal{F}_{p}$. Formally, the training objective $\mathcal{L}_{\text{{OS-Kairos}}}$ of the next-word prediction can be expressed as:
$\mathcal{L}_{\text{{OS-Kairos}}}$ $=\sum_{i=1}^{N}\mathcal{P}_{\theta}((a_{t}||\text{score}_{t})^{i}\mid P_{p}(s_{t},$ (3)
$\tau_{i},h_{t-1},(a_{t}||\text{score}_{t})^{<i})),$
where $N$ is the token number of $a_{t}$ and $\text{score}_{t}$, ||||| | is the concatenated operator of the prediction of action and score, and $\theta$ is the trainable parameters in OS-Kairos. This optimization is more stable compared to multi-task learning, as it not only preserves OS-Kairos’s action prediction ability but also generates confidence in the predicted actions.
Models API SCROLL PRESS STOP CLICK TYPE Total TSR
Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$
GPT-4o 22.22 100.00 46.67 86.95 74.63 93.62 90.07 87.59 76.35 39.13
GLM-4V-Plus 0.00 0.00 20.00 95.88 37.65 21.99 20.57 81.57 33.80 4.35
Qwen-VL-MAX 0.00 100.00 92.21 51.25 38.33 96.45 92.21 58.73 46.89 29.81
Auto-UI 44.44 0.00 0.00 2.93 0.15 0.00 0.00 2.81 0.59 0.00
Qwen2-VL-7B 22.22 85.71 0.00 37.98 15.69 55.32 42.55 40.75 20.49 0.00
OS-Atlas-Pro-7B 66.67 0.00 20.00 97.80 62.46 99.29 63.12 95.90 61.36 14.29
OS-Kairos 100.0033.33↑ 100.00100.00↑ 100.0080.00↑ 99.852.05↑ 96.3333.87↑ 100.000.71↑ 92.8629.74↑ 99.883.98↑ 95.9034.54↑ 88.2073.91↑
Table 2: Comparison of OS-Kairos with baselines in complex scenarios (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate relative improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
Adaptive Interaction GUI Agent. To ensure interactive adaptivity, we introduce a threshold to control OS-Kairos’s sensitivity. Formally, for a given threshold $\gamma$, OS-Kairos satisfies:
$f_{\text{confidence}}:\langle a_{t},\text{score}_{t}\rangle\xrightarrow{<\gamma}\text{Interactive},$ (4)
where human intervention is triggered if the current action’s confidence is below $\gamma$, otherwise, it is automatic. Notably, OS-Kairos switches to autonomous mode if the $\gamma$ is set to minimum value, or to fully interactive GUI if it is set to maximum value.
5 Experiments
This section will introduce the experimental setup, followed by our empirical results and analysis.
5.1 Experiment Setup
Datasets. Thanks to the confidence probing framework, we can evaluate the OS-Kairos in complex scenarios by splitting the generated trajectories. Moreover, we evaluate it on established benchmarks such as AITZ and Meta-GUI. Details are provided in Appendix C.2.
Models. In the confidence probing framework, we use the open-source GUI agent OS-Atlas-Pro-7B Wu et al. (2024b) as our probing model. Our objective is to probe the confidence score of the GUI model at each step, thereby introducing OS-Kairos to enhance its effectiveness. Additionally, we use $\mathtt{GPT}$-$\mathtt{4o}$ Achiam et al. (2023) as our critic model. The layout-parse model is resnet18 and convnextTiny for OCR-detection and recognition models, respectively Tang et al. (2019).
Baselines. We compare the proposed OS-Kairos with the following types:
$\bullet$ Multimodal API-based models. We consider MLLM-powered GUI agents, including $\mathtt{GPT}$-$\mathtt{4o}$ Achiam et al. (2023), $\mathtt{GLM}$-$\mathtt{4V}$-$\mathtt{Plus}$ GLM et al. (2024) and $\mathtt{Qwen}$-$\mathtt{VL}$-$\mathtt{MAX}$ Bai et al. (2023), which are strong baselines in zero-shot settings.
$\bullet$ Multimodal Open-source models. In the zero-shot setting, we also consider GUI-adapted MLLMs, including CogAgent Hong et al. (2024), Auto-UI Zhang and Zhang (2024), Qwen2-VL-7B Bai et al. (2023), OS-Atlas-Pro-7B Wu et al. (2024b). In the fine-tuning setting, we compare OS-Kairos with fine-tuned models on datasets.
Metrics. Following Wu et al. (2024b), we report the action type accuracy (Type), step-wise success rate (SR), and task success rate (TSR). Besides, we evaluate the human intervention success rate (HSR), intervention precision (IP), autonomous precision (AP), and relative efficiency (RE). More details of metrics and implementation can be found in Appendix C.3 and Appendix C.4.
Models Mode SCROLL PRESS STOP CLICK TYPE Total TSR
Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$
OS-Kairos Dataset
Auto-UI FT 0.0044.44↓ 71.4371.43↑ 80.0080.00↑ 98.8395.90↑ 67.1667.01↑ 97.1697.16↑ 0.710.71↑ 96.9694.15↑ 55.7455.15↑ 2.482.48↑
Qwen2-VL-7B FT 55.5633.34↑ 42.8643.85↓ 100.00100.00↑ 98.8360.85↑ 85.3469.65↑ 99.2943.97↑ 90.7848.23↑ 98.4857.73↑ 85.8365.34↑ 62.1162.11↑
OS-Atlas-Pro-7B FT 22.2244.45↓ 14.2914.29↑ 93.3373.33↑ 99.561.76↑ 84.7522.29↑ 99.290.00↑ 91.4928.37↑ 98.712.81↑ 84.7823.42↑ 55.9041.61↑
OS-Kairos ZS 100.0033.33↑ 100.00100.00↑ 100.0080.00↑ 99.852.05↑ 96.3333.87↑ 100.000.71↑ 92.8629.74↑ 99.883.98↑ 95.9034.54↑ 88.2073.91↑
AITZ Benchmark
CogAgent FT 70.2213.81↑ 45.952.35↓ 24.6019.84↑ 88.238.33↑ 66.1514.64↑ 45.8021.60↓ 21.8010.20↓ 72.596.73↑ 53.288.76↓ /
Auto-UI FT 61.4013.48↓ 57.708.61↑ 74.4014.28↑ 74.5630.19↑ 32.2019.48↑ 87.8014.80↑ 81.4013.60↑ 82.989.19↑ 47.6913.23↑ /
Qwen2-VL-7B FT 71.3852.74↑ 21.850.66↑ 78.5778.57↑ 88.3017.25↑ 51.1018.21↑ 87.805.00↑ 45.000.00↑ 85.1418.86↑ 55.2326.98↑ 1.781.78↑
OS-Atlas-Pro-7B FT 62.2334.83↑ 28.4827.82↑ 73.6168.45↑ 90.752.56↓ 58.7423.87↑ 89.003.80↑ 44.0016.60↑ 86.691.49↑ 58.3224.66↑ 11.1511.15↑
OS-Kairos ZS 91.1763.77↑ 73.5172.85↑ 91.6586.49↑ 98.435.12↑ 89.4654.59↑ 99.2014.00↑ 72.8045.40↑ 96.8111.61↑ 87.5453.88↑ 24.5124.51↑
Meta-GUI Benchmark
Auto-UI FT 42.9517.95↓ 65.9165.91↑ 53.0853.08↑ 84.2357.33↑ 53.9951.30↑ 86.5586.55↑ 1.751.75↑ 73.0253.00↑ 48.4942.04↑ 20.4220.42↑
Qwen2-VL-7B FT 89.1089.10↑ 72.7372.73↑ 90.0289.59↑ 94.6143.07↑ 83.1980.86↑ 97.0859.07↑ 64.3346.20↑ 93.1757.27↑ 83.4380.39↑ 57.2957.08↑
OS-Atlas-Pro-7B FT 84.6268.59↑ 70.4570.45↑ 89.3889.38↑ 96.011.48↑ 85.5348.24↑ 95.9135.68↑ 65.5050.30↑ 93.4927.40↑ 84.2760.68↑ 57.2956.78↑
OS-Kairos ZS 99.3683.33↑ 100.00100.00↑ 94.7394.73↑ 99.815.28↑ 96.6659.37↑ 98.8338.60↑ 95.3280.12↑ 98.4932.40↑ 96.3672.77↑ 87.7187.29↑
Table 3: Comparison of OS-Kairos in the fine-tuning setting. ZS and FT denote zero-shot and fine-tuning evaluations, respectively. We report overall accuracy for Type, SR, and TSR, as well as fine-grained accuracy for each action. Subscripts indicate relative improvement over the ZS baseline, with the best result highlighted in bold.
5.2 Main Results
We present comparison results for complex scenarios and two benchmarks with zero-shot settings in Table  3, Appendix  C.6.1, and Appendix C.6.2. Table2 provides a comprehensive comparison of fine-tuning settings. Our key findings are as follows: In zero-shot setting: Superior Performance and Better Effectiveness. Without changing the model capabilities, OS-Kairos significantly outperforms the zero-shot baseline across three datasets, highlighting its effectiveness. Specifically, the adaptive interaction of OS-Kairos effectively identifies complex steps that trigger human intervention. This not only improves the prediction accuracy for each action but also enhances overall performance. For example, it achieves 95.90% in SR and 88.20% in TSR for complex scenarios. Although API-based and proprietary models realizes domain enhancement for GUI tasks, they cannot identify complex steps, resulting in over-execution and task failure. Moreover, OS-Kairos yields promising results on the other two datasets when applying confidence scoring integration to the original dataset, highlighting its generality (see Appendix C.6.3).
In fine-tuning setting: Competitive Performance and Precise Improvement. Although fine-tuning can alleviate the over-execution of GUI agents, OS-Kairos still outperforms them, achieving high SR and notable improvements in TSR. For example, it shows relative improvements ranging from 26.09% to 85.72% in complex scenarios. Furthermore, OS-Kairos achieves precise improvements by identifying complex steps (e.g., $\mathtt{SCROLL}$), while fine-tuning may introduce side effects in specific actions and encounter optimization bottlenecks.
Figure 4: Analysis of intervention precision.
Confidence Scoring: Effective interaction. In Figure 4, OS-Kairos shows accurate confidence evaluation (HSR). Thus, It does not interfere with the autonomous steps (AP), as seen in complex scenarios (96.44%) and MetaGUI (93.18%). Notably, OS-Kairos achieves over 70% precision in all human intervention steps, highlighting its effective interaction. With high-quality sampling, we consider OS-Kairos’s precision can be improved further (e.g., AITZ). Additionally, the ablation study of the critic model shows that $\mathtt{GPT}$-$\mathtt{4o}$ is the optimal choice (see Appendix C.7).
5.3 Analysis
5.3.1 Dynamic Evaluation of TSR
[…]
Models Human Steps Actual Steps RE (%$\uparrow$ TSR (%$\uparrow$
GPT-4o 229 302 75.83 36.00
Qwen2-VL-7B 229 397 57.68 4.00
OS-Atlas-Pro-7B 229 359 63.79 26.00
[…]
Table 4: Analysis of efficiency and dynamic TSR.
5.3.2 Efficiency Evaluation
Table 4 reports the efficiency in a real-world environment. First, we count the optimal number of human steps on 50 instructions, about 429 steps. Next, we evaluate the actual step counts for baseline and OS-Kairos, respectively. Notably, the model max steps are set to 10. We observe that the baseline models tend to over-execute when faced with a complex step. In contrast, OS-Kairos more closely resembles human manipulation of a GUI, achieving 86.42% and 93.47% in RE.
5.3.3 Comparing Prompt-based Interaction
Table 5 presents a comparison of OS-Kairos with prompt-based interactive models. We see that the interactive mechanism of OS-Kairos outperforms the prompt-based paradigm, particularly surpassing the prompt-based OS-Atlas-Pro-7B in terms of HSR. Despite the strong grounding capabilities of $\mathtt{GPT}$-$\mathtt{4o}$ and $\mathtt{GLM}$-$\mathtt{4V}$-$\mathtt{Plus}$, API-based agents present instability, resulting in over-execution and sub-optimal performance. Among open-source GUI agents, Qwen2-VL-7B performs more consistently than OS-Atlas-Pro-7B, because prompt-based interactive severely disrupts the latter’s instruction-following ability.
Models Interactive Type (%$\uparrow$ SR (%$\uparrow$ TSR (%$\uparrow$ HSR (%$\uparrow$
GPT-4o Prompt 88.80 79.25 46.58 /
GLM-4V-Plus Prompt 88.34 79.03 47.83 /
Qwen2-VL-7B Prompt 76.42 38.44 25.47 /
OS-Atlas-Pro-7B Prompt 59.02 95.67 9.94 0.00
OS-Kairos FT 99.88 95.90 88.20 86.87
Table 5: Analysis of interactive paradigms vs. prompt-based baseline in complex scenarios.
Figure 5: Generality of OS-Kairos across model scale.
5.4 General Effectiveness across Scales
Model Scale.
Although the dynamic detection framework is built on the OS-Atlas-Pro-7B backbone, confidence scores and actions generated are supposed to be downwardly compatible. In other words, weaker models can be enhanced through data distillation and confidence scoring integration. Figure 5 shows that OS-Kairos can be successfully generalized to the 2Bsimilar-$\sim$7B model. First, Type and SR are effective, guaranteeing a TSR of 76.40% on the Qwen2-VL-7B model, 77.64% on OS-Atlas-Pro-4B, and 85.09% on Qwen2-VL-2B. Thus, the combination of confidence scoring and data distillation will enhance weak models, thus satisfying the deployment in resource-constrained environments.
Data Scale.
To evaluate the effect of data scaling on confidence scoring integration, we divide the trajectories from the probing framework into different scales for training and test data. As shown in Table 6, OS-Kairos remain stable in Type and SR scores across scales. Benefiting from its high HSR, OS-Kairos’s TSR accuracy reaches 76.19$\sim$88.20%, proving that the integration of confidence scoring into OS-Kairos requires only a small number of probing data at a significantly lower cost than fine-tuning the GUI agent.
Data Scaling Type (%$\uparrow$ SR (%$\uparrow$ TSR (%$\uparrow$ HSR (%$\uparrow$
9:1 99.25 92.21 76.19 84.67
8:2 99.88 95.90 88.20 86.87
7:3 99.46 94.16 83.94 84.79
6:4 99.41 94.05 78.30 84.47
Table 6: Varying data scale in confidence scoring.
Figure 6: Threshold impact on interactive sensitivity.
5.4.1 Interactive Sensitivity
OS-Kairos use a threshold to achieve adaptive interaction. To analyze the adaptive interaction sensitivity of OS-Kairos, we ablate threshold $\gamma$ from 0 to 5. In Figure 6, TSR and SR increase with the rise in interactive sensitivity, indicating that human intervention enhances the effectiveness of GUI agents in complex scenarios. The HSR and Type accuracy remain stable across different thresholds, indicating that OS-Kairos can effectively identify complex steps, especially in coordinates and input scenarios, alleviating over-execution of the GUI agent. The ablation study of adaptive interaction shows that OS-Kairos is more flexible (See Appendix C.8).
6 Conclusion
This study identifies a key challenge of over-execution in GUI agents, which poses substantial risks in complex scenarios, such as those involving ambiguous user instructions, unexpected interruptions, and environmental hijacks. To address the challenge, we introduce OS-Kairos, an adaptive GUI agent capable of predicting confidence levels at each step and efficiently deciding whether to act autonomously or seek human intervention. Concretely, we propose a collaborative probing framework for annotating confidence scores at each interaction step. By integrating confidence scoring, OS-Kairos outperforms previous GUI agents and API-based models, with improved effectiveness, scalability, generality, and efficiency.
Limitation
We acknowledge two primary limitations in our study. First, we only sampled instructions from three typical complex scenarios, as our focus was to investigate why existing GUI agents struggle with TSR and generate action confidence scores without loss of generality. Notably, we demonstrate the effectiveness of OS-Kairos on the AITZ and Meta-GUI benchmarks, which provide additional diverse instructions for complex scenarios. Besides, the generalization capabilities of OS-Kairos can mitigate these limitations. Second, experiments were focused on our probing dataset and two benchmark datasets, highlighting the need for complex scene probing and confidence scoring integration. Given that confidence scoring relies on proprietary models and high-quality human sampling, we anticipate that future research will explore the optimization of our approach to confidence scoring and evaluate new benchmark datasets.
Ethics Statement
This section presents the ethics statements in the following aspects: (i) Privacy. The probing instructions are sourced from publicly available datasets, human designers, and $\mathtt{GPT4o}$, covering 12 apps and 12 topics. Temporary accounts were used to register these apps, and the trajectories generated by our collaborative probing framework are available, ensuring that no personal data or personally identifiable information was collected. The two benchmarks employed also implemented safeguards to protect privacy Zhang et al. (2024b); Sun et al. (2022). Moreover, OS-Kairos, as an open-source GUI agent that does not rely on any external information and supports local deployment. (ii) System security. OS-Kairos follows the first principles thinking (Zhang and Zhang, 2024), manipulates the GUI like a human being, and can initiate human intervention in scenarios involving system security to ensure safety. (iii) Potential social impacts. OS-Kairos can improve the effectiveness of GUI execution instructions. Unlike fully autonomous GUI agents, OS-Kairos will proactively request authorization and acquire personal information, thus reducing malicious abuse.
Appendix A Why GUI agents have poor TSR?
In our pilot experiment, the TSR of autonomous GUI agents is significantly lower than interactive GUI agents. This difference is attributed to the poor exact match for SR, particularly for the $\mathtt{CLICK}$ and $\mathtt{TYPE}$ actions. In contrast, interactive GUI agents can mitigate this limitation through human intervention. Intuitively, we consider the impact of exact matching on trajectory steps to be exponential. Formally, for a trajectory with $k$ steps, the probability that instruction $\tau_{i}$ can be completed is:
$\text{{TSR}}_{\tau_{i}}=\prod_{j=1}^{k}\text{{SR}}_{j},s.t.,\textit{SR}_{j}\sim\beta(u,l).$ (5)
[…]
$\ln(\textit{TSR}_{{\tau_{i}}})\sim\mathcal{N}\left(k\cdot\mu\right.\left.,\,k\cdot\sigma^{2}\right),$ (6)
then,
$\textit{TSR}_{{\tau_{i}}}$ $\sim\text{LogNormal}(\exp^{k\mu+\frac{k\sigma^{2}}{2}},\,$ (7)
$\exp^{2k\mu+k\sigma^{2}}(\exp^{k\sigma^{2}}-1)).$
[…]
When generalized to $N$ independent and identically distributed instructions, the average TSR satisfies:
$\text{{TSR}}_{avg}=\frac{1}{N}\sum_{i=1}^{N}\prod_{j=1}^{k}\text{{SR}}_{j}.$ (8)
According to center limit theory, TSRavg also satisfies normal distribution:
$\textit{TSR}_{{avg}}\sim\mathcal{N}(\exp^{k\mu+\frac{k\sigma^{2}}{2}},\,$ (9)
$\exp^{2k\mu+k\sigma^{2}}(\exp^{k\sigma^{2}}-1)/N).$
As shown in Figure 7(c), we observe the exponential effect of SR on TSR. In the autonomous mode, TSRavg,${}_{\text{avg},\text{auto}}$ is nearly 0%. In contrast, OS-Kairos and the fully interactive GUI agent both achieve success rates exceeding 60%.
[…]
$0\approx m^{\delta}\cdot q^{k-\delta}<\text{{TSR}}_{\tau_{i}}<m^{\delta}\cdot p^{k-\delta}\approx p^{k}.$ (10)
Figure 8: Illustration of the effect of SR on TSRavg for $N$ trajectories.
[…]
Appendix B Instruction Distribution
In our dynamic capability probing, we collect 1,000 instructions for three complex scenarios, covering 12 topics and 13 apps. The distributions of topics and apps are shown in Figure 9 and Figure 10. The distribution of scenarios is shown in Figure 11.
Figure 9: Subject distribution of instructions.
Figure 10: APP distribution of instructions.
Figure 11: Distribution of steps in different scenarios.
Appendix C Implementation Details
C.1 Prompt Templates
Below are the prompt templates for designing OS-Kairos. In the collaborative probing framework, we design a planning prompt, action phase prompt, action prompt for probed GUI agent $\mathcal{F}_{p}$ and critic model $\mathcal{F}_{c}$, scoring prompt, and finishing judgment prompt. In the confidence-driven interaction phase, we only use the $\mathcal{F}_{p}$ action prompt to optimize and evaluate OS-Kairos. The pipeline controller fills these variables in them according to the context.
C.1.1 Planning Prompt Template
In the GUI capability probing framework, the critic model $\mathcal{F}_{c}$ generates the planning schedule of the user instruction based on the $\mathtt{GPT}$-$\mathtt{4o}$ and the instruction planning prompt, as shown in Figure 12.
Figure 12: Prompts of the critic model for generating the planning schedule of the user instruction.
C.1.2 Action Phase Prompt Template
In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ determines the current step based on the $\mathtt{GPT}$-$\mathtt{4o}$ and action phase prompts, as shown in Figure 13.
Figure 13: Prompts of the critic model for generating the action phase.
C.1.3 Action Prompt Template
In the collaborative probing framework, we obtain the prediction of GUI agents using action prompts. The action prompts for the probed GUI agent $\mathcal{F}_{p}$ and critic model $\mathcal{F}_{c}$ are shown in Figure 14 and Figure 15, respectively. Following Wu et al. (2024b); Zhang et al. (2024b), we define the actions set, which comprises 7 kinds of actions: $\mathtt{CLICK}$, $\mathtt{SCROLL}$, $\mathtt{TYPE}$, $\mathtt{PRESS\_BACK}$, $\mathtt{PRESS\_HOME}$, $\mathtt{COMPLETE}$, and $\mathtt{IMPOSSIBLE}$.
Figure 14: Prompt of the probed GUI agent for generating action.
Figure 15: Prompt of the critic model for generating action.
C.1.4 Scoring Prompt Template
In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ generates the score for the action of $\mathcal{F}_{p}$ based on $\mathtt{GPT}$-$\mathtt{4o}$ and scoring prompt, as shown in Figure 16.
Figure 16: Prompt of the critic model for generating action score.
C.1.5 Completion Judgment Prompt Template
In the collaborative probing framework, the critic model $\mathcal{F}_{c}$ exploits $\mathtt{GPT}$-$\mathtt{4o}$ to judge whether the instruction is completed. The prompt is shown in Figure 17.
Figure 17: Prompt for the critic model to judge instruction completion.
C.2 Details of Datasets
We consider evaluating OS-Kairos on three customized complex scenarios and two benchmarks: AITZ Zhang et al. (2024b) and Meta-GUI Sun et al. (2022). The statistics of the dataset are shown in the Table 7.
  • AITZ Zhang et al. (2024b): The first dataset to employ chain-of-action thought (CoAT) connects perception (of screen layouts and UI elements) with cognition (of action decision-making) to enhance the AITW benchmark. This dataset comprises 2,504 operation trajectories across 18.6K real-world intentions. Based on the application domain, AITZ is also divided into five subsets: General, Install, GoogleApps, Single, and WebShopping.
  • Meta-GUI Sun et al. (2022): task-oriented dialogue dataset is released for interactive GUI agent. These utterances cut a trajectory into several dialogue turns. Meta-GUI consists of 1,009 trajectories with 16.4K steps. The data diversity lies in 11 applications of 6 topics.
OS-Kairos Trajectory Screen Goal
Train 800 4078 759
Test 200 1054 198
AITZ Trajectory Screen Goal
General 479 3607 479
Install 420 3627 420
Google Apps 242 1889 242
Single 844 2594 844
Web Shopping 519 6926 519
Meta-GUI Trajectory Screen Goal
Train 897 14539 2286
Test 116 1923 336
Table 7: Dataset statistics.
C.3 Details of Evaluation Metrics
To ensure fair comparison across all baseline methods, we standardize the evaluation metrics for each action. We define the SR metrics for the three complex actions as follows:
  • $\mathtt{CLICK}$: GUI agent predictions are considered correct if and only if both action types and position coordinates <x, y>. Following Zhang and Zhang (2024), we measure performance by calculating the distance between the predicted and ground truth coordinates. We consider the coordinates to be correct if the distance between the coordinates and the ground truth is within 14% of the screen width.
  • $\mathtt{TYPE}$: GUI agent predictions are considered correct if and only if both action type and action content are correct.
  • $\mathtt{SCROLL}$: GUI agent predictions are considered correct if and only if both action type and direction argument (i.e., UP, DOWN, LEFT, and RIGHT) are correct.
Furthermore, Type measures the exact match score between the predicted action types (e.g., CLICK, SCROLL) and the ground truth. TSR requires that all steps in a trajectory be correctly executed. For HSR, we define four statistical metrics with the threshold $\gamma$:
  • True positive (TP): Neither the prediction confidence nor the ground truth exceeds the $\gamma$, i.e., the agent does not require and perform interactions.
  • False positive (FP): The prediction confidence is greater than the $\gamma$, but the ground truth does not, meaning the agent is not required, but interaction is performed.
  • True negative (TN): Both the prediction confidence and the ground truth exceed the $\gamma$, meaning the agent must also perform interaction.
  • False negative (FN): The prediction confidence is less than $\gamma$, but the ground truth is greater than $\gamma$, which means that the agent needs but does not perform interactions.
Hence, HSR can be calculated:
$\textit{HSR}=\frac{TP+TN}{TP+TN+FP+FN}.$ (11)
In addition, IP calculates the accuracy of the intervention step where intervention is actually needed, while AP measures the accuracy of the autonomous step where autonomy is truly required. Hence, IP and AP can be calculated:
$\textit{IP}=\frac{TN}{TN+FN},AP=\frac{TP}{TP+FP}$ (12)
Following Wang et al. (2024a), RE measures the relative efficiency of the GUI agent compared to the steps taken by humans. It demonstrates whether OS-Kairos can use the mobile device more efficiently.
C.4 Implement Details
For each dataset, we randomly split 80% trajectories as training data, and 20% trajectories as testing data. Dataset statistics are presented in Table 7 of Appendix C.2. To ensure a fair comparison with the baseline, we use $\mathtt{GPT}$-$\mathtt{4o}$ to score between the probing model and ground truth actions in the two benchmarks, without relying on high-quality sampling. In the zero-shot scenario, we evaluate the GUI agent directly using prompt learning. In the fine-tuning scenario, we fine-tune the model for 8 epochs on the corresponding dataset with a learning rate of 1e-5. In the interactive mode, if not specifically mentioned, the threshold $\gamma$ is set to 4. When human intervention is required at the current step, OS-Kairos uses ground truth for the evaluation of the data set or human guidance for the dynamic evaluation. Our experiments are conducted on 8$\times$×NVIDIA A100 80 GB GPUs.
C.5 Usage of Existing Artifacts
For API-based MLLMs, we access them directly via the official interface. For open-source MLLMs, we either download the model weights from Hugging Face [^1] https://huggingface.co/models or reproduce the model using the same training strategy. In our proposed OS-Kairos, the layout-parse pipeline of the collaborative probing framework is built upon Modelscope  [^2] https://modelscope.cn/home. Furthermore, we utilize $\mathtt{LLaMA}$-$\mathtt{Factory}$ [^3] https://github.com/hiyouga/LLaMA-Factory to fine-tune the probed model on three datasets for confidence integration. Notably, the InternVL-based models are fine-tuned using $\mathtt{Xtuner}$ [^4] https://github.com/InternLM/xtuner. All licenses for these packages permit their use for standard academic research purposes.
C.6 Further Analysis
C.6.1 AITZ Benchmark
Table 8 presents a comparison of OS-Kairos with the baselines in the AITZ benchmark. In API-based MLLMs, although $\mathtt{GPT}$-$\mathtt{4o}$ performs the best, it is nearly impossible to finish user instructions. Among the open-source GUI agents, OS-Atlas-Pro-7B outperforms the other baselines due to the adaptation of AITZ, but still exhibits low SR and cannot fully complete user instructions. In contrast, OS-Kairos achieves precise intervention in complex steps on top of OS-Atlas-Pro-7B, with significant improvements in actions and overall performance. As a result, OS-Kairos’s TSR increased from 0% to 24.51%.
Models API SCROLL PRESS STOP CLICK TYPE Total TSR
Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$
GPT-4o 24.17 23.84 0.00 63.80 27.71 35.20 16.00 58.32 22.69 0.00
GLM-4V-Plus 11.65 7.28 0.00 79.15 27.65 43.80 20.40 68.95 20.92 0.00
Qwen-VL-MAX 7.89 13.04 10.2 / 72.3 / 34.04 / 52.41 /
CogAgent 56.41 48.30 4.76 79.90 51.50 67.40 34.00 65.86 44.52 /
Auto-UI 74.88 49.09 60.12 44.37 12.72 73.00 67.80 73.79 34.46 /
Qwen2-VL-7B 18.64 21.19 0.00 71.05 32.89 82.80 45.00 66.28 28.25 0.00
OS-Atlas-Pro-7B 27.40 0.66 5.16 93.31 34.87 85.20 27.40 85.20 33.66 0.00
OS-Kairos 91.1763.77↑ 73.5172.85↑ 91.6586.49↑ 98.435.12↑ 89.4654.59↑ 99.2014.00↑ 72.8045.40↑ 96.8111.61↑ 87.5453.88↑ 24.5124.51↑
Table 8: Comparison of OS-Kairos with baselines in the AITZ benchmark (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate relative improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
C.6.2 Meta-GUI Benchmark
Meta-GUI benchmark dataset is an out-of-domain (OOD) task against probing models, which allows for probing more complex steps and generating the confidence level for each step. Table 9 presents the performance of OS-Kairos on the Meta-GUI benchmark compared to the baseline. First, API-based MLLMs exhibit lower SR (17.19% to 32.72%) and Type (54.74% to 69.85%), which can be attributed to over-execution on complex steps such as $\mathtt{SCROLL}$. Hence, Qwen-VL-MAX only achieves a TSR of 15.42%, while $\mathtt{GLM}$-$\mathtt{4v}$-$\mathtt{Plus}$ performs weakly, with only 1.67% TSR. In addition, three open-source GUI agents such as OS-Atlas-Pro-7B are even less effective, as they cannot adapt to OOD instructions. In contrast, OS-Kairos achieves the accuracies of 98.49% in Type, 96.36% in SR and 87.71% in TSR, respectively. Similarly, the fine-grained Type and SR outperform the baseline.
Models API SCROLL PRESS STOP CLICK TYPE Total TSR
Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$ Type (%) $\uparrow$ SR (%) $\uparrow$
GPT-4o 33.97 25.00 12.79 94.12 42.30 66.47 28.14 69.85 32.72 6.67
GLM-4V-Plus 0.00 0.00 1.06 95.45 26.53 38.01 22.81 65.05 17.19 1.67
Qwen-VL-MAX 14.74 40.91 1.91 70.87 37.85 74.85 45.03 54.74 27.86 15.42
Auto-UI 60.90 0.00 0.00 26.90 2.69 0.00 0.00 20.02 6.45 0.00
Qwen2-VL-7B 0.00 0.00 0.43 51.54 2.33 38.01 18.13 35.90 3.04 0.21
OS-Atlas-Pro-7B 16.03 0.00 0.00 94.53 37.29 60.23 15.20 66.09 23.59 0.42
OS-Kairos 99.3683.33↑ 100.00100.00↑ 94.7394.73↑ 99.815.28↑ 96.6659.37↑ 98.8338.60↑ 95.3280.12↑ 98.4932.40↑ 96.3672.77↑ 87.7187.29↑
Table 9: Comparison of OS-Kairos with baselines in the Meta-GUI benchmark (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate relative improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
C.6.3 Generality Evaluation of OS-Kairos
Figure 18: Generality analysis of OS-Kairos for adaptive interaction from original dataset to target datasets.
OS-Kairos outperforms the baseline model across three datasets due to the integration of the confidence scoring. To verify the generality of adaptive interaction, we train OS-Kairos on each of the three datasets and then test it on the other two. The evaluation results are presented in Figure 18. We see that OS-Kairos is able to achieve a decent performance, though the domains vary. Compared to the main results, it significantly outperforms the baseline in the zero-shot setting across three datasets, particularly in SR and TSR metrics. Also, its generalization performance is comparable to that of models fine-tuned directly on the target dataset. We also note that the more complex the dataset on which confidence scoring is integrated, the better the generalization of OS-Kairos. For example, OS-Kairos exhibits the best generalization with confidence scoring integration on the Meta-GUI dataset (29.05% vs. 21.74% in the AITZ benchmark, and 83.85% vs. 88.20% in the OS-Kairos dataset).
C.7 Ablation of Critic Models
As the advanced judgment capabilities of $\mathtt{GPT}$-$\mathtt{4o}$ Chen et al. (2024a), we utilize it as the critic model in the collaborative probing framework. To analyze the impact of the critic model on OS-Kairos confidence integration and GUI adaptive interaction, we select Qwen-VL-MAX as an alternative. Table 10 presents the adaptive interaction performance of OS-Kairos across different critic models. The results show that the scoring quality of $\mathtt{GPT}$-$\mathtt{4o}$ significantly outperforms Qwen-VL-Max, with an HSR of 86.87% compared to 57.63%. In addition, the precision of the intervention decreases by 4.59% in the autonomous steps and 9.25% in the complex steps. Although GUI performance is similar, Qwen-VL-Max leads to more frequent interventions with OS-Kairos.
Models Type (%$\uparrow$ SR (%$\uparrow$ TSR (%$\uparrow$ HSR (%$\uparrow$ IP(%$\uparrow$ AP(%$\uparrow$
GPT-4o 98.49 96.36 87.71 86.87 70.75 96.44
Qwen-VL-MAX 99.65 96.01 85.71 57.63 61.50 91.55
Table 10: Ablation of critic models.
C.8 Ablation of Adaptive Interaction
To understand the advantages of adaptive integration in OS-Kairos, we compare its performance with and without adaptive integration: when treated as regression optimization or classification optimization in complex scenarios datasets. As shown in Table 11, we see that OS-Kairos without adaptive interaction is quite accurate at HSR of 86.99%, but its overall performance and intervention precision are suboptimal. For example, the TSR is 82.61%, IP is 70.66%, and AP is 95.84%. The results show that the adaptive interaction does not significantly affect the performance of OS-Kairos. In contrast, OS-Kairos has the advantage of adaptive interaction by tuning the threshold, which balances the sensitivity between autonomous and human intervention.
Models Type (%$\uparrow$ SR (%$\uparrow$ TSR (%$\uparrow$ HSR (%$\uparrow$ IP(%$\uparrow$ AP(%$\uparrow$
OS-Kairos 99.88 95.90 88.20 86.87 70.75 96.44
OS-Kairosw/o adaptive ${}_{\text{w/o adaptive interaction}}$ 99.53 95.31 82.61 86.99 70.66 95.84
Table 11: Ablation of adaptive interaction.
Appendix D Case study
To further illustrate the execution process of OS-Kairos, we present four examples from three complex scenarios, along with two examples from the benchmark datasets. First, for simple instructions, OS-Kairos can be fully autonomous, as shown in Figure 19. Second, for complex instructions across the three scenarios, OS-Kairos adaptively identifies the complex steps requiring human intervention, while automating other steps, as shown in Figure 20, Figure 21 and Figure 22. Similarly, OS-Kairos performs effectively on the AITZ benchmark (Figure 23). In an extreme case, OS-Kairos requests human intervention at nearly every step to complete the task, as the Meta-GUI benchmark represents an OOD scenario for OS-Atlas-Pro-7B, as shown in Figure 24.
Figure 19: Case study of OS-Kairos in the normal scenario. At each step, OS-Kairos outputs both the action and the confidence score. If the score falls below a specified threshold, human intervention is initiated to ensure task success.
Figure 20: Case study of OS-Kairos in Scenario 1 (capability bottleneck). At each step, OS-Kairos outputs both the action and the confidence score. If the score falls below a specified threshold, human intervention is initiated to ensure task success.
Figure 21: Case study of OS-Kairos in Scenario 2 (no location permission). At each step, OS-Kairos outputs both the action and the confidence score. If the score falls below a specified threshold, human intervention is initiated to ensure task success.
Figure 22: Case study of OS-Kairos in Scenario 3 (information absence). At each step, OS-Kairos outputs both the action and the confidence score. If the score falls below a specified threshold, human intervention is initiated to ensure task success.
Figure 23: Case study of OS-Kairos in the AITZ benchmark. At each step, OS-Kairos outputs both the action and the confidence score. If the score falls below a specified threshold, human intervention is triggered to ensure task success.
Figure 24: Case study of OS-Kairos in the Meta-GUI benchmark. At each step, OS-Kairos outputs both the action and the confidence score. If the score falls below a specified threshold, human intervention is triggered to ensure task success.